the cost of hiring a car for 2 days in 2006 was 264$ which was 20% more than what it was in 2001.what was the cost of hiring a car in 2001 for 2 days.
Answer by Theo(13342)

x = cost in 2001.

formula is:

264 = 1.2 * x

solve for x to get x = 264 / 1.2 = 220.

cost in 2001 was 220.

cost in 2006 was 264.

(264 - 220) / 220 = .2 = 20% more cost in 2006 than in 2001.